How do i prove tan(theta)+cot(theta)/sec(theta)csc(theta)=1?
1
Expert's answer
2012-11-05T10:04:57-0500
sec(theta)=1/cos(theta)
csc(theta)=1/sin(theta)
tan(theta)=sin(theta)/cos(theta)
cot(theta)=cos(theta)/sin(theta)

( tan(theta)+cot(theta) )/( sec(theta)csc(theta))=[sin(theta)/cos(theta)+cos(theta)/sin(theta)]/[1/( sin(theta) * cos(theta) ) ]=
=[ (sin^2(theta)+cos^2 (theta) )/( sin(theta)cos(theta) )
]*[sin(theta)*cos(theta) ]=sin^2(theta)+cos^2 (theta)=1